circular flower garden that has a radius of 6.25 ft and is 9 inches deep. Need to buy topsoil to fill the garden which one bad cover 3 cubic feet and cost $3.38 per bag
Answer by rothauserc(4718)

circular flower garden that has a radius of 6.25 ft
:
depth of circular flower garden is 9/12 = 3/4 ft
:
Volume is pi * r^2 * depth = (22/7) * 6.25 * (3/4) = 14.7321 cubic ft
:
14.7321/3 = 4.9107 is approximately 5 bags
:
5 * 3.38 = 16.9
:
*************************************************
need to buy 5 bags of topsoil at a cost of $16.90
*************************************************
: